# Turing Video-based Cognitive Tests to Handle Entangled Concepts

### v1

@@ 3 A novel video-based cognitive test
Table 1: We report the statistical data collected in the video-based cognitive test presented in Section 3. The judgement probabilities are in substantial agreement with the results obtained in other text-based cognitive tests, namely, [30], [38] and [39], presented in Table 2. Also in this case, we get a significant violation of the CHSH inequalitiy in Equation (1), with a CHSH factor (see Equation (2)) exceeding Cirel’son’s bound.
@@
Table 2: We report in comparison the statistical data of three text-based cognitive tests, namely, the Aerts & Sozzo (2011) test [30] in the first column, the Aerts et al. (2023) test [38] in the second column, and the Bertini et al. (2023) test [39] in the third column. The corresponding judgement probabilities are in substantial agreement across the tests and also with the video-based cognitive test in Table 1. All tests exhibit a significant violation of the CHSH inequality.
4 Data modelling in Hilbert space
In this section, we provide a quantum representation in Hilbert space of the data collected in the video-based cognitive test in Section 3. We apply the general quantum theoretical framework that we have elaborated for the modelling of any Bell-type test, as we have already done in various papers [31, 34, 37, 38].
The quantum theoretical framework for Bell-type situations consists in the implementation of three main steps, as follows.
(ii) Whenever an individual participating in the test reads the introductory text which explains the details of the test and nature of the concepts involved, this set of instructions prepares the composite entity The Animal Acts in an initial state $p$ which describes the general situation of an animal that produces a recognizable sound. Each participant is then confronted with this uniquely prepared state $p$. More precisely, in the coincidence measurement $XY$, $X=A,A^{\prime}$, $Y=B,B^{\prime}$, each participant interacts with the entity The Animal Acts in the state $p$ and operates as a measurement context for the entity. This interaction generally changes, in an intrinsically indeterministic way, $p$ into a new state depending on the choice that is made, as a consequence of this ‘contextual interaction’. E.g., if the participant chooses in $AB$ The Horse Whinnies, which corresponds to the outcome $A_{1}B_{2}$ (see Section 3), the interaction between the entity The Animal Acts in the state $p$ and the (mind of the) participant determines an indeterministic change of state of the entity from $p$ to the state $p_{A_{1}B_{2}}$ which describes the more concrete situation of a horse that whinnies. More generally, for every $X=A,A^{\prime}$, $Y=B,B^{\prime}$, the coincidence measurement $XY$ has four possible outcomes $X_{i}Y_{j}$, where we choose $X_{i}Y_{j}=+1$ if $i=j$ and $X_{i}Y_{j}=-1$ if $i\neq j$, and four outcome states, or eigenstates, $p_{X_{i}Y_{j}}$, describing the state of The Animal Acts after the outcome $X_{i}Y_{j}$ occurs in $XY$, $i,j=1,2$. When all responses are collected, a statistics of the outcomes $X_{i}Y_{j}$ arises which is interpreted, in the large number limit, as the probability $P_{p}(X_{i}Y_{j})$ that the outcome $X_{i}Y_{j}$ is obtained when the coincidence measurement $XY$ is performed on the composite entity The Animal Acts in the initial state $p$.
(iii) We have identified the entities, initial state, coincidence measurements, outcome probabilities and eigenstates that are relevant to the The Animal Acts situation. Then, we have to work out a quantum representation in Hilbert space to model the data collected in this situation, that is, the entity The Animal Acts is associated with a complex Hilbert space and the initial state $p$ is represented by a unit vector of this Hilbert space. Next, for every $X=A,A^{\prime}$, $Y=B,B^{\prime}$, the coincidence measurement $XY$ is represented by a self-adjoint operator or, equivalently, by a spectral family, on the Hilbert space whose eigenvectors represent the eigenstates of $XY$, while the outcome probabilities are obtained from Born’s rule of quantum probability.
Regarding the above Hilbert space representation, we preliminarily observe that all coincidence measurements $XY$, $X=A,A^{\prime}$, $Y=B,B^{\prime}$, have four outcomes $X_{i}Y_{j}$, $i,j=1,2$, which entails that the composite entity The Animal Acts is associated, as an overall entity, with the complex Hilbert space $\mathbb{C}^{4}$ of all ordered 4-tuples of complex numbers. Moreover, each state $p$ of The Animal Acts is represented by a unit vector of $\mathbb{C}^{4}$ and each coincidence measurement on The Animal Acts is represented by a self-adjoint operator or, equivalently, by a spectral family, on $\mathbb{C}^{4}$. On the other hand, for every $i,j=1,2$, each outcome $X_{i}Y_{j}$ is obtained by juxtaposing the outcomes $X_{i}$ and $Y_{j}$, e.g., The Tiger Growls, is obtained by syntactically juxtaposing the words “tiger” and “growls”. This operation defines a 2-outcome measurement $X$, $X=A,A^{\prime}$, on the individual entity Animal and a 2-outcome measurement $Y$, $Y=B,B^{\prime}$, on the individual entity Acts. Hence, each of these individual entities is associated with the complex Hilbert space $\mathbb{C}^{2}$ of all ordered couples of complex numbers. Should we had performed separate measurements on Animal and Acts, the Hilbert space formalism would have prescribed that the composite entity The Animal Acts would have been associated with the tensor product $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$. But, we remind that we are studying here the identification problem (see Section 2), that is, the problem of how the composite entity The Animal Acts can be decomposed into the individual entities Animal and Acts in such a way that these individual entities can be recognised from measurements performed on the composite entity. As such, we are doing an operation that is the inverse of what one typically does in Bell-type situations in quantum physics, where one constructs or, better, composes, the measurements on the composite entity from measurements performed on individual entities.
@@ 4 Data modelling in Hilbert space
With reference to a Bell-type setting, one can now prove that, if the coincidence measurements ${XY}$ and ${XY^{\prime}}$, $X=A,A^{\prime}$, $Y,Y^{\prime}=B,B^{\prime}$, $Y\neq Y^{\prime}$, are product measurements with respect to the isomorphism $I$, then, for every state $p$ of the composed entity, the marginal law of Kolmogorovian probability expressed by Equation (3) is satisfied. Analogously, if the coincidence measurements $XY$ and $X^{\prime}Y$, $X,X^{\prime}=A,A^{\prime}$, $Y=B,B^{\prime}$, $X\neq X^{\prime}$, are product measurements with respect to the isomorphism $I$, then, for every state $p$ of the composed entity, the marginal law expressed by Equation (4) is satisfied. One also proves that, if the marginal law is satisfied in all coincidence measurements, then a unique isomorphism exists, which can be chosen to be the identity operator [31].
It follows from the above that, if the marginal law is violated, then one cannot find a unique isomorphism between $\mathbb{C}^{4}$ and $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ such that all measurements are product measurements with respect to this isomorphism. In this case, one cannot explain the violation of the CHSH inequality as due to the usual situation in quantum physics where all measurements are product measurements and only the initial pre-measurement state is entangled
 [^6] We add that there are reasons to believe that the marginal law is also violated in typical Bell-type tests on quantum physical entities, which indicates that entangled measurements are involved also in physical domain. However, the experimental violation of the marginal law in these tests is not large, hence has hardly been reflected about from a theoretical point of view
. Furthermore, if the marginal law is systematically violated, as it occurs in our test (see Table 1), then four distinct isomorphisms $I_{XY}$, exist such that the measurement ${XY}$ is a product measurement with respect to $I_{XY}$, $X=A,A^{\prime}$, $Y=B,B^{\prime}$. As a consequence, there is no unique isomorphism allowing to identify individual entities of a given composite entity. Finally, if we consider a given isomorphism between $\mathbb{C}^{4}$ and $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ with respect to which identifying individual entities of a composite entity in a given test, then it may happen that both the initial pre-measurement state and all measurements are entangled [31].
@@
Firstly, the non-classical correlations that violate the CHSH inequality in Equation (1) in the The Animal Acts situation can be reasonably attributed to the fact that ‘the component concepts carry meaning and further meaning is created in the combination process’. Since the violation of the CHSH inequality indicates the presence of entanglement between the individual conceptual entities, then it is reasonable that ‘it is the quantum structure of entanglement that theoretically capture the meaning that is non-classically created in this case’. This suggests that the initial state $p$ of the composite entity The Animal Acts should be an entangled state.
Secondly, in the The Animal Acts situation, since all coincidence measurements ${XY}$, $X=A,A^{\prime}$, $Y=B,B^{\prime}$, violate the marginal law of Kolmogorovian probability, also these measurements should be entangled measurements. In addition, in each measurement $XY$, all outcomes $X_{i}Y_{j}$, correspond to combined concepts, though less abstract than The Animal Acts, e.g., in The Cat Meows, meaning is created with respect to Cat and Meows taken separately, which suggests that all eigenstates $p_{X_{i}Y_{j}}$, $i,j=1,2$, should also be entangled states.
@@
which represents the maximally entangled state corresponding to the singlet spin state, as typically done in Bell-type tests in quantum physics (see Section 2). However, before making this choice here too, it is worth to reflect about the general modelling scheme we adopted to represent the states of conceptual entities (see, e.g., [25]). In [52] we used the unit vector
@@
where $\mu(A_{i}B_{j})$ are the judgement probabilities of the coincidence measurement $AB$ and the unit vectors $|p_{A_{i}B_{j}}\rangle$ form an ON basis of eigenvectors representing the eigenstates of $AB$, $i,j=1,2$, in the 4-dimensional Hilbert space describing the $AB$-measurement situation. This representation is in agreement with the general modelling scheme in [25], because the unit vector in Equation (12) represents the initial state of The Animal Acts in a measurement having The Horse Growls, The Horse Whinnies, The Bear Growls, and The Bear Whinnies as possible outcomes. Analogously, one could use, with obvious changes of symbols, the unit vectors
@@
to represent the initial state of The Animal Acts when the coincidence measurement $AB^{\prime}$ is performed, the unit vector
@@
to represent the initial state of The Animal Acts when the coincidence measurement $A^{\prime}B$ is performed, and the unit vector
@@
to represent the initial state of The Animal Acts when the coincidence measurement $A^{\prime}B^{\prime}$ is performed. Indeed, also the unit vectors in Equations (13), (14) and (15) reproduce the correct judgement probabilities in the corresponding coincidence measurements.
How will we then represent the initial state of The Animal Acts with respect to the overall experiment that tests the CHSH version of Bell’s inequalities according to the general modelling scheme in [25]? The straightforward answer is to take the normalized superposition state represented by the linear combination
@@
Let us recall that the considered basis vectors for each of the coincidence measurements are not necessarily the same, their relation depending on how these coincidence measurements relate experimentally to the CHSH form of the Bell’s test experiment, and only experiments to test these relations can give us this exact information. This however does not avoid the linear combination in Equation (16) to be well defined and representing the initial state of The Animal Acts with respect to the CHSH form of the Bell’s test.
@@
The considerations above indicate that the unit vector in Equation (16) is the most appropriate candidate to represent the initial state corresponding to a preparation of the entity The Animal Acts according to the general modelling scheme in [25]. However, for the experiments where the data are obtained by calculating relative frequencies of occurrence of different choices, using separately the fours vectors in Equations (12), (13), (14) and (15) comes to the same concerning the predictions made by the model, because relative frequencies of occurrence of outcomes are always determined separately for each of the four coincidence cases. Or, more concretely, for these experiments we have, e.g.,
@@
Since the outcomes in cognitive tests with human participants are to a large extent also determined by estimating relative frequencies of occurrence, one can expect that in this case the same, this time almost, equalities. Indeed, since some human participants in the tests will have their choices determined in a totally different way, these equalities do not hold in this case, although even those participants will have the tendency not to let their choice about one of the coincidence pairs depend on that they gave answers before or after for the other coincidence pairs, which again makes the effect on outcomes for the two states rather little different, and hence we can speak of ‘almost equalities’. Hence, given the complexity of the human mind, it is not excluded that the answers given for one of the coincidence pairs are indeed partly determined by the presence in the same test of the other coincidence pairs, which means that for psychology experiments we should in principle consider both states giving rise to different predictions for the to be rested probabilities.
@@
The above is a justifiable choice if we think of a preparation that is very minimal and leads to the test person considering a kind of bare The Animal Acts. For example, before the instruction sheet is read by the test person, it could have been communicated that it involves measurements involving the sentence “the animal acts”. For this minimal initial preparation, that maximally entangled state could serve in the quantum theoretical model we are building also according to the general modelling scheme in [25].
We wish to digress briefly in connection with this problem of precisely determining the prepared state, firstly and foremost with the intention of fully clarifying it, and secondly because an element emerges here where our approach shows its strength in connection with a possible use for AI based on quantum structures. Even if we wish to determine the prepared state of an experiment where only the sentence “the animal acts” is presented to the test subjects, the singlet spin state is not the best choice to serve as the prepared state. A superposition state in which all animals and all actions that can be performed figure would be the real correct mathematical representation. This would lead however to having to consider a Hilbert space of giant dimension as tensor product of all these possibilities. The human mind is probably not capable of handling this condition, but does make an attempt in that direction. Studies on how concepts are represented in the human mind preferentially point to working with, on the one hand, a precedence for the representation of some leading exemplars (see, e.g., [55] on exemplar theories of concepts) and, on the other hand, with the representation of a prototype of the concept in question (see, e.g., [56] on prototype theories of concepts). In any case, if these exemplar and prototype theories are correct, they also point to a shortcoming of the human mind in connection with such a representation, and it can be expected that a powerful AI would indeed be able to represent such a giant superposition in every detail. Indeed, in the end, the collection of all animals and all possible actions that each is capable of and that can be conceptually described is still a finite collection.
@@
Then, coming to measurements, let us represent the coincidence measurements ${XY}$, $X=A,A^{\prime}$, $Y=B,B^{\prime}$. As we have seen above, each measurement has four outcomes $X_{i}Y_{j}=\pm 1$ and four eigenstates $p_{X_{i}Y_{j}}$, $i,j=1,2$. For every $X=A,A^{\prime}$, $Y=B,B^{\prime}$, we represent $XY$ by the spectral family defined by the ON basis of the four eigenvectors $|p_{X_{i}Y_{j}}\rangle$, where we set, for every $i,j=1,2$,
@@
This corresponds to four conditions for each coincidence measurement $XY$.
@@
This corresponds to six additional conditions for each coincidence measurement $XY$.
@@
where we have used Born’s rule in Equation (19). This corresponds to four additional conditions for each coincidence measurement $XY$.

### v2

@@ 4 A novel video-based cognitive test
Table 1: We report the statistical data collected in the video-based cognitive test presented in Section 4. The judgement probabilities are hun in substantial agreement with the results obtained in other text-based cognitive tests, namely, Aerts & Sozzo, (2011), Aerts et al., 2023b and Bertini Lemporini & Moriani, (2023), presented in Table 2. Also in this case, we get a significant violation of the CHSH inequalitiy in Equation (1), with a CHSH factor (see Equation (2)) exceeding Cirel’son’s bound.
5 Data modelling in Hilbert space
In this section, we provide a quantum representation in Hilbert space of the data collected in the video-based cognitive test in Section 4. We apply the general quantum theoretical framework that we have elaborated for the modelling of any Bell-type test, as we have already done in various articles (Aerts & Sozzo,, 2014; Aerts et al.,, 2019, 2021; Aerts et al., 2023b, ) . The quantum theoretical framework for Bell-type situations consists in the implementation of three main steps, as follows.
(ii) Whenever a person participating in the test reads the introductory text which explains the details of the test and nature of the concepts involved, this set of instructions prepares the composite entity The Animal Acts in an initial state $p$ which describes the general situation of an animal that produces a recognizable sound. Each participant is then confronted with this uniquely prepared state $p$. More precisely, in the joint measurement $XY$, $X=A,A^{\prime}$, $Y=B,B^{\prime}$, each participant interacts with the entity The Animal Acts in the state $p$ and operates as a measurement context for the entity. This interaction generally changes, in an intrinsically indeterministic way, $p$ into a new state depending on the choice that is made, as a consequence of this ‘contextual interaction’. E.g., if the participant chooses in $AB$ The Horse Whinnies, which corresponds to the outcome $A_{1}B_{2}$ (see Section 4), the interaction between the entity The Animal Acts in the state $p$ and the (mind of the) participant determines an indeterministic change of state of the entity from $p$ to the state $p_{A_{1}B_{2}}$ which describes the more concrete situation of a horse that whinnies. More generally, for every $X=A,A^{\prime}$, $Y=B,B^{\prime}$, the joint measurement $XY$ has four possible outcomes $X_{i}Y_{j}$, where we choose $X_{i}Y_{j}=+1$ if $i=j$ and $X_{i}Y_{j}=-1$ if $i\neq j$, and four outcome states, or eigenstates, $p_{X_{i}Y_{j}}$, describing the state of The Animal Acts after the outcome $X_{i}Y_{j}$ occurs in $XY$, $i,j=1,2$. When all responses are collected, a statistics of the outcomes $X_{i}Y_{j}$ arises which is interpreted, in the large number limit, as the probability $P_{p}(X_{i}Y_{j})$ that the outcome $X_{i}Y_{j}$ is obtained when the joint measurement $XY$ is performed on the composite entity The Animal Acts in the initial state $p$.
(iii) We have identified the entities, initial state, joint measurements, outcome probabilities and eigenstates that are relevant to the The Animal Acts situation. Then, we have to work out a quantum representation in Hilbert space to model the data collected in this situation, that is, the entity The Animal Acts is associated with a complex Hilbert space and the initial state $p$ is represented by a unit vector of this Hilbert space. Next, for every $X=A,A^{\prime}$, $Y=B,B^{\prime}$, the joint measurement $XY$ is represented by a self-adjoint operator or, equivalently, by a spectral family, on the Hilbert space whose eigenvectors represent the eigenstates of $XY$, while the outcome probabilities are obtained from Born’s rule of quantum probability.
Regarding the Hilbert space representation above, we preliminarily observe that all joint measurements $XY$, $X=A,A^{\prime}$, $Y=B,B^{\prime}$, have four outcomes $X_{i}Y_{j}$, $i,j=1,2$, which entails that the composite entity The Animal Acts is associated, as an overall entity, with the complex Hilbert space $\mathbb{C}^{4}$ of all ordered 4-tuples of complex numbers. Moreover, each state $p$ of The Animal Acts is represented by a unit vector of $\mathbb{C}^{4}$ and each joint measurement on The Animal Acts is represented by a self-adjoint operator or, equivalently, by a spectral family, on $\mathbb{C}^{4}$. On the other hand, for every $i,j=1,2$, each outcome $X_{i}Y_{j}$ is obtained by juxtaposing the outcomes $X_{i}$ and $Y_{j}$, e.g., The Tiger Growls, is obtained by syntactically juxtaposing the words “tiger” and “growls”. This operation defines a 2-outcome measurement $X$, $X=A,A^{\prime}$, on the individual entity Animal and a 2-outcome measurement $Y$, $Y=B,B^{\prime}$, on the individual entity Acts. Hence, each of these individual entities is associated with the complex Hilbert space $\mathbb{C}^{2}$ of all ordered pairs of complex numbers. Should we had performed separate measurements on Animal and Acts, the Hilbert space formalism would have prescribed that the composite entity The Animal Acts would have been associated with the tensor product $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$. But, we remind that we are studying here the identification problem (see Section 2), that is, the problem of how the composite entity The Animal Acts can be decomposed into the individual entities Animal and Acts in such a way that these individual entities can be recognised from measurements performed on the composite entity. As such, we are doing an operation that is the inverse of what one typically does in Bell-type situations in quantum physics, where one constructs or, better, composes, the measurements on the composite entity from measurements performed on individual entities.
@@ 5 Data modelling in Hilbert space
Table 2: We report in comparison the statistical data of three text-based cognitive tests, namely, the Aerts & Sozzo (2011) test (Aerts & Sozzo,, 2011) in the first column, the Aerts et al. (2023) test (Aerts et al., 2023b, ) in the second column, and the Bertini et al. (2023) test (Bertini Lemporini & Moriani,, 2023) in the third column. The corresponding judgement probabilities are in substantial agreement across the tests and also with the video-based cognitive test in Table 1. All tests exhibit a significant violation of the CHSH inequality.
@@
With reference to a Bell-type setting, one can now prove that, if the joint measurements ${XY}$ and ${XY^{\prime}}$, $X=A,A^{\prime}$, $Y,Y^{\prime}=B,B^{\prime}$, $Y\neq Y^{\prime}$, are product measurements with respect to the isomorphism $I$, then, for every state $p$ of the composed entity, the marginal law condition expressed by Equation (3) is satisfied. Analogously, if the joint measurements $XY$ and $X^{\prime}Y$, $X,X^{\prime}=A,A^{\prime}$, $Y=B,B^{\prime}$, $X\neq X^{\prime}$, are product measurements with respect to the isomorphism $I$, then, for every state $p$ of the composed entity, the marginal law condition expressed by Equation (4) is satisfied. One also proves that, if the marginal law conditions are satisfied in all joint measurements, then a unique isomorphism exists, which can be chosen to be the identity operator (Aerts & Sozzo,, 2014).
It follows from the above that, if the marginal law conditions are violated, then one cannot find a unique isomorphism between $\mathbb{C}^{4}$ and $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ such that all measurements are product measurements with respect to this isomorphism. In this case, one cannot explain the violation of the CHSH inequality as due to the usual situation in quantum physics where all measurements are product measurements and only the initial pre-measurement state is entangled. [^7] As anticipated in Section 2, there are reasons to believe that the marginal law conditions are also violated in typical Bell-type tests on quantum physical entities, which indicates that entangled measurements are involved also in the physical domain. However, the violation of the marginal law conditions in these tests is not large, hence has hardly been reflected about from a theoretical point of view (Aerts et al.,, 2019). Furthermore, if the marginal law conditions are systematically violated, as it occurs in our test (see Table 1), then four distinct isomorphisms $I_{XY}$, exist such that the measurement ${XY}$ is a product measurement with respect to $I_{XY}$, $X=A,A^{\prime}$, $Y=B,B^{\prime}$. As a consequence, there is no unique isomorphism allowing to identify individual entities of a given composite entity. Finally, if we consider a given isomorphism between $\mathbb{C}^{4}$ and $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ with respect to which identifying individual entities of a composite entity in a given test, then it may happen that both the initial pre-measurement state and all measurements are entangled (Aerts & Sozzo,, 2014).
@@
Firstly, the non-classical correlations that violate the CHSH inequality in Equation (1) in the The Animal Acts situation can be reasonably attributed to the fact that ‘the component concepts carry meaning and further meaning is created in the combination process’. Since the violation of the CHSH inequality indicates the presence of entanglement between the individual conceptual entities, then it is reasonable that ‘it is the quantum structure of entanglement that theoretically captures the meaning that is non-classically created in this case’. This suggests that the initial state $p$ of the composite entity The Animal Acts should be an entangled state.
Secondly, in the The Animal Acts situation, since all joint measurements ${XY}$, $X=A,A^{\prime}$, $Y=B,B^{\prime}$, violate the marginal law conditions of Kolmogorovian probability, also these measurements should be entangled measurements. In addition, in each measurement $XY$, all outcomes $X_{i}Y_{j}$, correspond to combined concepts, though less abstract than The Animal Acts, e.g., in The Cat Meows, meaning is created with respect to Cat and Meows taken separately, which suggests that all eigenstates $p_{X_{i}Y_{j}}$, $i,j=1,2$, should also be entangled states.
@@
which represents the maximally entangled state corresponding to the singlet spin state, as typically done in Bell-type tests in quantum physics (see Section 2). However, before making this choice here too, it is worth to reflect about the general modelling scheme we adopted to represent the states of conceptual entities (see, e.g., Aerts et al., (2021)). In Aerts & Beltran, (2022) we used the unit vector
@@
where $\mu(A_{i}B_{j})$ are the judgement probabilities of the joint measurement $AB$ and the unit vectors $|p_{A_{i}B_{j}}\rangle$ form an ON basis of eigenvectors representing the eigenstates of $AB$, $i,j=1,2$, in the 4-dimensional Hilbert space describing the $AB$-measurement situation. This representation is in agreement with the general modelling scheme in Aerts et al., (2021), because the unit vector in Equation (12) represents the initial state of The Animal Acts in a measurement having The Horse Growls, The Horse Whinnies, The Bear Growls, and The Bear Whinnies as possible outcomes. Analogously, one could use, with obvious changes of symbols, the unit vectors
@@
to represent the initial state of The Animal Acts when the joint measurement $AB^{\prime}$ is performed, the unit vector
@@
to represent the initial state of The Animal Acts when the joint measurement $A^{\prime}B$ is performed, and the unit vector
@@
to represent the initial state of The Animal Acts when the joint measurement $A^{\prime}B^{\prime}$ is performed. Indeed, also the unit vectors in Equations (13), (14) and (15) reproduce the correct judgement probabilities in the corresponding joint measurements.
How will we then represent the initial state of The Animal Acts with respect to the overall experiment that tests the CHSH version of Bell’s inequalities according to the general modelling scheme in Aerts et al., (2021)? The straightforward answer is to take the normalized superposition state represented by the linear combination
@@
Let us recall that the considered basis vectors for each of the joint measurements are not necessarily the same, their relation depending on how these joint measurements relate experimentally to the CHSH form of the Bell’s test experiment, and only experiments to test these relations can give us this exact information. This however does not avoid the linear combination in Equation (16) to be well-defined and representing the initial state of The Animal Acts with respect to the CHSH form of the Bell’s test.
@@
The considerations above indicate that the unit vector in Equation (16) is the most appropriate candidate to represent the initial state corresponding to a preparation of the entity The Animal Acts according to the general modelling scheme in Aerts et al., (2021). However, for the experiments where the data are obtained by calculating relative frequencies of occurrence of different choices, using separately the fours vectors in Equations (12), (13), (14) and (15) comes to the same concerning the predictions made by the model, because relative frequencies of occurrence of outcomes are always determined separately for each of the four joint cases. Or, more concretely, for these experiments we have, e.g.,
@@
Since the outcomes in cognitive tests with human participants are to a large extent also determined by estimating relative frequencies of occurrence, one can expect that in this case the same, this time almost, equalities. Indeed, since some human participants in the tests will have their choices determined in a totally different way, these equalities do not hold in this case, although even those participants will have the tendency not to let their choice about one of the joint pairs depend on that they gave answers before or after for the other joint pairs, which again makes the effect on outcomes for the two states rather little different, and hence we can speak of ‘almost equalities’. Hence, given the complexity of the human mind, it is not excluded that the answers given for one of the joint pairs are indeed partly determined by the presence in the same test of the other joint pairs, which means that for psychology experiments we should in principle consider both states giving rise to different predictions for the to be rested probabilities.
@@
The above is a justifiable choice if we think of a preparation that is very minimal and leads to the test person considering a kind of bare The Animal Acts. For example, before the instruction sheet is read by the test person, it could have been communicated that it involves measurements involving the sentence “the animal acts”. For this minimal initial preparation, that maximally entangled state could serve in the quantum theoretical model we are building also according to the general modelling scheme in Aerts et al., (2021).
We wish to digress briefly in connection with this problem of precisely determining the prepared state, firstly and foremost with the intention of fully clarifying it, and secondly because an element emerges here where our approach shows its strength in connection with a possible use for AI based on quantum structures. Even if we wish to determine the prepared state of an experiment where only the sentence “the animal acts” is presented to the test subjects, the singlet spin state is not the best choice to serve as the prepared state. A superposition state in which all animals and all actions that can be performed figure would be the real correct mathematical representation. This would lead however to having to consider a Hilbert space of giant dimension as tensor product of all these possibilities. The human mind is probably not capable of handling this condition, but does make an attempt in that direction. Studies on how concepts are represented in the human mind preferentially point to working with, on the one hand, a precedence for the representation of some leading exemplars (see, e.g., Nosofski, (1988) on exemplar theories of concepts) and, on the other hand, with the representation of a prototype of the concept in question (see, e.g., Rosch, (1973) on prototype theories of concepts). In any case, if these exemplar and prototype theories are correct, they also point to a shortcoming of the human mind in connection with such a representation, and it can be expected that a powerful AI would indeed be able to represent such a giant superposition in every detail. Indeed, in the end, the collection of all animals and all possible actions that each is capable of and that can be conceptually described is still a finite collection.
@@
Then, coming to measurements, let us represent the joint measurements ${XY}$, $X=A,A^{\prime}$, $Y=B,B^{\prime}$. As we have seen above, each measurement has four outcomes $X_{i}Y_{j}=\pm 1$ and four eigenstates $p_{X_{i}Y_{j}}$, $i,j=1,2$. For every $X=A,A^{\prime}$, $Y=B,B^{\prime}$, we represent $XY$ by the spectral family defined by the ON basis of the four eigenvectors $|p_{X_{i}Y_{j}}\rangle$, where we set, for every $i,j=1,2$,
@@
This corresponds to four conditions for each joint measurement $XY$.
@@
This corresponds to six additional conditions for each joint measurement $XY$.
@@
where we have used Born’s rule in Equation (19). This corresponds to four additional conditions for each joint measurement $XY$.